OpenStudy (anonymous):

What is the 'lim x approaches 0 sin3x/7x'? With work please

OpenStudy (tkhunny):

Well, show your work, please. How else are you going to get it? Hint: \(\dfrac{\sin(3x)}{7x} = \dfrac{3}{7}\cdot\dfrac{\sin(3x)}{3x}\)